OpenStudy (anonymous):

among 10 laptop 5 are good and 5 have defects .unaware of this a customer buys 6 laptops (a) what is the probability that exactly two of them are defective (b)given that atleast 2 purchases laptops are defective,what is the probability that exactly 2 are defective?

OpenStudy (anonymous):

Take note of the formula for the binomial probability.

OpenStudy (anonymous):

\[\displaystyle P(x) = \frac{n!}{x!*(n-x)!}*p^x*(1-p)^{n-x}\]

OpenStudy (anonymous):

where n is the sample size, p is the probability of success, and x is the number of successes.

OpenStudy (anonymous):

The factorial works like 10! = 10*9*8*7*6*5*4*3*2*1 = 3628800..